# A Survey of Protocol Fuzzing

### v2

@@ Acknowledgements.
- Zhu et al. (2022a) Xiaogang Zhu, Sheng Wen, Seyit Camtepe, and Yang Xiang. 2022a. Fuzzing: A Survey for Roadmap. ACM Comput. Surv. (jan 2022). https://doi.org/10.1145/3512345 Just Accepted.
- Zhu et al. (2022b) Xiaogang Zhu, Sheng Wen, Seyit Camtepe, and Yang Xiang. 2022b. Fuzzing: A Survey for Roadmap. ACM Comput. Surv. (jan 2022). https://doi.org/10.1145/3512345 Just Accepted.
@@
- Zou et al. (2020) Qingtian Zou, Anoop Singhal, Xiaoyan Sun, and Peng Liu. 2020. Generating Comprehensive Data with Protocol Fuzzing for Applying Deep Learning to Detect Network Attacks. CoRR abs/2012.12743 (2020). arXiv:2012.12743 https://arxiv.org/abs/2012.12743
@@
- Zou et al. (2021) Yong-Hao Zou, Jia-Ju Bai, Jielong Zhou, Jianfeng Tan, Chenggang Qin, and Shi-Min Hu. 2021. TCP-Fuzz: Detecting Memory and Semantic Bugs in TCP Stacks with Fuzzing. In 2021 USENIX Annual Technical Conference (USENIX ATC 21). USENIX Association, 489–502. https://www.usenix.org/conference/atc21/presentation/zou
@@
- Zuo et al. (2022) Feilong Zuo, Zhengxiong Luo, Junze Yu, Ting Chen, Zichen Xu, Aiguo Cui, and Yu Jiang. 2022. Vulnerability Detection of ICS Protocols via Cross-State Fuzzing. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems 41, 11 (2022), 4457–4468. https://doi.org/10.1109/TCAD.2022.3201471
@@
- Zuo et al. (2021) Feilong Zuo, Zhengxiong Luo, Junze Yu, Zhe Liu, and Yu Jiang. 2021. PAVFuzz: State-Sensitive Fuzz Testing of Protocols in Autonomous Vehicles. 2021 58th ACM/IEEE Design Automation Conference (DAC) (2021), 823–828.

### v3

@@ Acknowledgements.
- Zhu et al. (2022) Xiaogang Zhu, Sheng Wen, Seyit Camtepe, and Yang Xiang. 2022. Fuzzing: A Survey for Roadmap. Comput. Surveys 54, 11s (2022).
@@
- Zou et al. (2020) Qingtian Zou, Anoop Singhal, Xiaoyan Sun, and Peng Liu. 2020. Generating Comprehensive Data with Protocol Fuzzing for Applying Deep Learning to Detect Network Attacks. (2020). arXiv:2012.12743
@@
- Zou et al. ([n. d.]) Yong-Hao Zou, Jia-Ju Bai, Jielong Zhou, Jianfeng Tan, Chenggang Qin, and Shi-Min Hu. [n. d.]. TCP-Fuzz: Detecting Memory and Semantic Bugs in TCP Stacks with Fuzzing. In 2021 USENIX Annual Technical Conference (USENIX ATC). 489–502.
@@
- Zuo et al. (2022) Feilong Zuo, Zhengxiong Luo, Junze Yu, Ting Chen, Zichen Xu, Aiguo Cui, and Yu Jiang. 2022. Vulnerability Detection of ICS Protocols via Cross-State Fuzzing. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems 41, 11 (2022), 4457–4468.
@@
- Zuo et al. (2021) Feilong Zuo, Zhengxiong Luo, Junze Yu, Zhe Liu, and Yu Jiang. 2021. PAVFuzz: State-Sensitive Fuzz Testing of Protocols in Autonomous Vehicles. 58th ACM/IEEE Design Automation Conference (DAC) (2021), 823–828.